OpenStudy (anonymous):

C sqrt(c) + 12c sqrt(c) - 2c sqrt(c) 13c sqrt(c) - 2c sqrt(c) 11c sqrt (c) <------------ Is that the right answer?

OpenStudy (anonymous):

is that first C the number of the question?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

your answer is correct i think

OpenStudy (anonymous):

\[c\sqrt c + 12c \sqrt c - 2c \sqrt c =11c\sqrt c\]\]

OpenStudy (anonymous):

yeah. thats the original problem. So i did it right?

OpenStudy (anonymous):

yes you did

OpenStudy (anonymous):

Yay! finally i understand Algebra . Thanks (:

OpenStudy (anonymous):

yw